OpenStudy (anonymous):

integral of e^x+1 / e^x -1 ?

OpenStudy (solomonzelman):

\[\LARGE\color{blue}{ \int\limits_{ }^{ } \frac{e^x+1}{e^x-1} }\] expand it to \[\LARGE\color{blue}{ \int\limits_{ }^{ } \frac{e^x}{e^x-1}~~dx+\int\limits_{ }^{ } \frac{1}{e^x-1}~~dx }\]

OpenStudy (anonymous):

i did that that thing....but im stucked in the 2nd equation

OpenStudy (solomonzelman):

\[\LARGE\color{blue}{ \int\limits_{ }^{ } \frac{1}{u}~~du+\int\limits_{ }^{ } \frac{1}{e^x-1}~~dx }\]

OpenStudy (solomonzelman):

Tell me what you get for the first integral

OpenStudy (anonymous):



OpenStudy (anonymous):

im thinking of multiplying it with\[\frac{ 1 }{ e ^{x}-1 }\frac{ e ^{x}+1 }{ e ^{x}+1 }\]

OpenStudy (anonymous):

i would go with partial fractions

OpenStudy (solomonzelman):

Tell me what you get for the first integral

OpenStudy (anonymous):

\[\frac{e^x+1}{e^x-1}=\frac{e^x-1+2}{e^x-1}=1+\frac{2}{e^x-1}\] first is obvious, second use \(u=e^x-1\)

OpenStudy (solomonzelman):

yes and then sub s for e^x (I think)

OpenStudy (anonymous):

i would go the whole hog with the \(u\) sub \[u=e^x-1\\ u+1=e^x\\ \log(u+1)=x\] making \[dx=\frac{1}{u+1}\] and the integral would then be \[2\int \frac{du}{u(u+1)}\]

OpenStudy (anonymous):

i got it now tnx